OpenStudy (anonymous):

the 24 people at the thanksgiving dinner, 1/8 of the people ate pumpkin pie. 1/2 ate cherry pie. 1/4 ate pecan pie. 1/8 ate apple pie. how many people ate pie?

OpenStudy (anonymous):

1/8 of 24 equals 3, so three people ate apple pie at Thanksgiving

OpenStudy (mathmate):

\(\frac{1}{8}+\frac{1}{8}+\frac{1}{4}+\frac{1}{2}=1\) and they all ate some kind of pie. So how many people ate pie?